# Generalized Models for Inflationary Preheating: Oscillations and Symmetries

### v2

@@ I Introduction
This paper considers a particular preheating scenario, where the inflaton field features a quartic potential during its oscillatory phase of preheating. The quartic potential results in nonlinear oscillations of the inflaton field. These nonlinearities, in turn, result in different forms for the variables that drive parametric resonance. More specifically, while parametric resonance is often described by the Mathieu equation Mathieu (1868), this scenario leads to a more general Hill’s equation Hill (1886); Magnus and Winkler (1966), which has different bands of instability (and stability) for resonance. More significantly, the effective forcing parameter in Hill’s equation grows with continued expansion of the universe, thus leading to more instability. This behavior stands in contrast to most previously considered cases where the evolution of the parameters in the Mathieu equation leads to greater stability. While the scenario of this paper considers the inflaton field to have a quartic form during the reheating epoch, the potential can have a more general form at earlier times (Özsoy et al., 2017). During the part of inflation when the universe expands exponentially, usually operating under slow-roll conditions, a quartic form for the potential is currently disfavored by observations of the cosmic microwave background. These observations constrain the form of the inflaton potential for times corresponding to 70 to 60 e-foldings before the end of the slow-roll era, whereas the reheating epoch takes place after its conclusion.
This paper is organized as follows. We first discuss (in Section II) how Effective Field Theory (EFT) can lead to different forms for the inflaton potential during the slow-roll epoch (when observable density fluctuations are produced) and the subsequent reheating epoch. Next we solve for the nonlinear oscillations of the inflaton field for the case of a quartic potential (in Section III), and derive the corresponding Hill’s equation for reheating. We then carry out a Floquet analysis of the resulting model equation and find its stability diagram (in Section IV). As outlined above, the expansion of the universe drives the forcing parameter in Hill’s equation to larger values, leading to greater instability. If the inflaton potential has a mass term as well as a quartic term during reheating, then the behavior of the reheating field is more complicated, and we address this issue in Section V. The paper concludes in Section VI with a summary of results and a discussion of their implications.
@@ II Challenges for Reheating and the EFT Approach
An alternative approach to understanding the dynamics of reheating is to use methods of Effective Field Theory (EFT) to describe the process Özsoy et al. (2017, 2015). Such methods have been useful in particle physics Petrov and Blechman (2016); Penco (2020), dark energy Park et al. (2010); Bloomfield et al. (2013); Gubitosi et al. (2013); Wen et al. (2022) and condensed matter physics Brauner et al. (2022). In the present context, the crucial feature of the EFT is that it allows us to establish universality classes Özsoy et al. (2017, 2015). The cosmic expansion for a given equation of state is determined, but the dynamics and possible departures can be classified in symmetry groups, which define the universality classes. In the case of dark energy, for example, it has been shown that different ways of triggering cosmic acceleration come in such classes respecting their common symmetries Bloomfield et al. (2013); Gubitosi et al. (2013); Wen et al. (2022). What we would like to emphasize in this paper is that symmetries could be a broader mechanism for discovery the dynamics of inflation reheating. The inflaton need not be a scalar field. In addition, it is important to emphasize the work of Greene and Kofman (2000) which first demonstrated that fermion preheating is also possible and efficient  [^1] We thank our journal referee for reminding of this important point and it indeed was part of our motivation for broadening our approach to the notion of (p)reheating to be more general than simple scalar fields and their interactions.. Instead, it can be a ‘clock’ or spurion of broken time-translation invariance. As we learn more about the nature of inflation this creates a useful framework, or methodology for exploring (p)reheating.
@@
This paper focuses on conformal symmetry and its subsequent breaking during the epochs following inflation [^2] In the case of conformal symmetry it is important to note that although we are taking a different approach to the reasoning there has been significant past related work, in particular Özsoy et al. (2015, 2017). The point here is that the EFT is a different way of thinking about the end of inflation and we have also tried to present further results beyond the initial use of this method.. Given that we do not know how inflation ends — and whether inflation is driven by a true scalar field — we approach this problem by considering the inflaton to be an order parameter, more specifically a clock that depends on time and but not on space. As a result, the inflaton breaks time-translation invariance (more precisely time diffeomorphism invariance), but it need not be a true scalar field. One can think of the field both as a clock and as a Goldstone boson of the broken time translation invariance. When considering cosmological solutions, here an FLRW universe, the cosmic expansion breaks the symmetry in time. This theoretical notion has (previously) led to the EFT description of inflation and can be continued for the case of reheating.
@@ III Reheating Model with Nonlinear Oscillations
then the leading order form of the potential will have a quartic form when $\phi$ is near its minimum. [^4] Note that one has to translate the field $\phi\to\phi-f$, in order to obtain the first term in equation [8]). Note that the potential (8) represents the simplest model of this type. In particular, the interaction term is linear in $\phi$ and quadratic in $\chi$, which results in a linear equation of motion of the reheating field and requires that the coupling field $\sigma$ is sufficiently small (so that the coupling term does not interfere with the oscillations of the inflaton in its quartic potential). In addition to having a different form at earlier epochs, the inflaton potential could have additional terms (e.g., $V\sim m^{2}\phi^{2}$) during reheating, as well as coupling to additional fields. These complications are considered in subsequent sections below. The interaction between $\phi$ and $\chi$ may take other forms as well, and conformal interaction terms are of particular interest considering the previous discussion. We show in Appendix B that such couplings don’t give favorable results to leading order, so we consider the trilinear interaction throughout this exploration of the quartic inflaton potential’s effects on reheating.
@@ IV Parametric Resonance
Figure 3: Left: Sample trajectories in $(A,q)$ space for $m_{\chi}/\omega_{0}=0$ (top row) and $m_{\chi}/\omega_{0}=2$ (bottom row) for trilinear quartic preheating. The black $A=2q$ line roughly separates the unstable region above from the stable region below. Right: corresponding growth rates $\gamma$ along trajectories as a function of scale factor $a$. We can see that a massless reheat field grows indefinitely without nonlinear effects, while a massive reheat field eventually stops growing on its own when the trajectory reaches a more stable area. Colors correspond to comoving wavenumbers $k/\omega_{0}$ chosen to result in the various $A_{0}$ shown, and the initial inflaton amplitude $\sqrt{\lambda}\Phi_{0}/\omega_{0}$ is chosen to result in $q_{0}\approx 50$.
@@ V Inclusion of a Mass Term in the Inflaton Potential
Figure 5: Growth rates $\gamma$ along trajectories as a function of scale factor $a$ corresponding to the trajectories in Figure 4. The solid lines correspond to the mixed case $m\neq 0,\ \lambda\neq 0$, the dashed lines to purely quartic $m=0,\ \lambda\neq 0$, and the dotted lines to purely massive $m=0,\ \lambda\neq 0$. The three columns correspond from left to right to $\lambda\Phi_{0}^{2}/m^{2}=1,10,1000$ in the mixed potential, and the parameters for the massive and quartic cases are picked to match $(A_{0},q_{0})$ of the mixed cases. The colors correspond to different wavenumbers $k_{0}/\omega_{0}$; note that for the massive case, $\omega_{0}=m$, for the quartic case, $\omega_{0}\approx 0.85\sqrt{\lambda}\Phi_{0}$, and for the mixed case, $\omega$ is the function shown in equation (54). We can see that mass of either the inflaton or reheat field result in an eventual decrease in the growth rate.
@@
VI Conclusion
@@ A.2 Optimization of the Correction Parameter
Appendix B Exclusion of quartic terms
@@
B.1 Quartic reheat field potential
A term $V(\chi)\supset\frac{1}{4}\lambda_{\chi}\chi^{4}$ adds a strong restorative force that suppresses the exponential increase in the reheat field’s amplitude, even with low strength $\lambda_{\chi}<1$. The trilinear equation of motion becomes, with the familiar definitions of $A$ and $q$,
@@ B.1 Quartic reheat field potential
The coefficient increases, $A\to A_{\mathrm{eff}}\equiv A+\lambda_{\chi}\chi_{k}^{2}$, which stabilizes the growth in $\chi_{k}$. The corresponding growth rates are shown in Figure 7.
The reheating potential and the inflation potential may arise from different EFTs and contain different dominant terms, so we do not require $V(\chi)\supset\frac{1}{4}\lambda_{\chi}\chi^{4}$ to bound the potential from below Dufaux et al. (2006). The inflaton itself is bounded because it oscillates during reheating, so we do not need to worry about the possibility of $V(\phi,\chi)\supset\sigma\phi\chi^{2}\to-\infty$.
@@
B.2 Forced oscillator solution for first quartic coupling
@@
B.3 Subdominant growth from second quartic coupling

### v3

@@ I Introduction
This paper considers a particular preheating scenario, where the inflaton field features a quartic potential during its oscillatory phase of preheating, building on the seminal work of Greene et al. (1997). The quartic potential results in nonlinear oscillations of the inflaton field. These nonlinearities, in turn, result in different forms for the variables that drive parametric resonance. More specifically, while parametric resonance is often described by the Mathieu equation Mathieu (1868), this scenario leads to a more general Hill’s equation Hill (1886); Magnus and Winkler (1966), which has different bands of instability (and stability) for resonance. More significantly, the effective forcing parameter in Hill’s equation grows with continued expansion of the universe, thus leading to more instability. This behavior stands in contrast to most previously considered cases where the evolution of the parameters in the Mathieu equation leads to greater stability. We build upon the methods of Greene et al. (1997) and show the Floquet maps for a wide range of scenarios with a primary focus on the efficient trilinear interaction, and determine the trajectories of the system through parameter space. In addition, we include numerical results to support our semianalytic work.
While the scenario of this paper considers the inflaton field to have a quartic form during the reheating epoch, the potential can have a more general form at earlier times (Özsoy et al., 2017). During the part of inflation when the universe expands exponentially, usually operating under slow-roll conditions, a quartic form for the potential is currently disfavored by observations of the cosmic microwave background. These observations constrain the form of the inflaton potential for times corresponding to 70 to 60 e-foldings before the end of the slow-roll era, whereas the reheating epoch takes place after its conclusion.
This paper is organized as follows. We first discuss (in Section II) how Effective Field Theory (EFT) can lead to different forms for the inflaton potential during the slow-roll epoch (when observable density fluctuations are produced) and the subsequent reheating epoch. Next we solve for the nonlinear oscillations of the inflaton field for the case of a quartic potential (in Section III), and derive the corresponding Hill’s equation for reheating. We then carry out a Floquet analysis of the resulting model equation and find its stability diagram (in Section IV). As outlined above, the expansion of the universe drives the forcing parameter in Hill’s equation to larger values, leading to greater instability. If the inflaton potential has a mass term as well as a quartic term during reheating, then the behavior of the reheating field is more complicated, and we address this issue in Section V. The paper concludes in Section VII with a summary of results and a discussion of their implications.
@@ II Challenges for Reheating and the EFT Approach
An alternative approach to understanding the dynamics of reheating is to use methods of Effective Field Theory (EFT) to describe the process Özsoy et al. (2017, 2015). Such methods have been useful in particle physics Petrov and Blechman (2016); Penco (2020), dark energy Park et al. (2010); Bloomfield et al. (2013); Gubitosi et al. (2013); Wen et al. (2022) and condensed matter physics Brauner et al. (2022). In the present context, the crucial feature of the EFT is that it allows us to establish universality classes Özsoy et al. (2017, 2015). The cosmic expansion for a given equation of state is determined, but the dynamics and possible departures can be classified in symmetry groups, which define the universality classes. In the case of dark energy, for example, it has been shown that different ways of triggering cosmic acceleration come in such classes respecting their common symmetries Bloomfield et al. (2013); Gubitosi et al. (2013); Wen et al. (2022). What we would like to emphasize in this paper is that symmetries could be a broader mechanism for discovery the dynamics of inflation reheating. The inflaton need not be a scalar field. In addition, it is important to emphasize the work of Greene and Kofman (2000) which first demonstrated that fermion preheating is also possible and efficient [^1] We thank our journal referee for reminding of this important point, and it indeed was part of our motivation for broadening our approach to the notion of (p)reheating to be more general than simple scalar fields and their interactions.. Instead, it can be a ‘clock’ or spurion of broken time-translation invariance. As we learn more about the nature of inflation this creates a useful framework, or methodology for exploring (p)reheating.
@@
This paper focuses on conformal symmetry and its subsequent breaking during the epochs following inflation [^2] In the case of conformal symmetry it is important to note that although we are taking a different approach to the reasoning, there has been significant past related work, in particular Özsoy et al. (2015, 2017). The point here is that the EFT is a different way of thinking about the end of inflation, and we have also tried to present further results beyond the initial use of this method.. Given that we do not know how inflation ends—and whether inflation is driven by a true scalar field—we approach this problem by considering the inflaton to be an order parameter, more specifically a clock that depends on time and but not on space. As a result, the inflaton breaks time-translation invariance (more precisely time diffeomorphism invariance), but it need not be a true scalar field. One can think of the field both as a clock and as a Goldstone boson of the broken time translation invariance. When considering cosmological solutions, here an FLRW universe, the cosmic expansion breaks the symmetry in time. This theoretical notion has (previously) led to the EFT description of inflation and can be continued for the case of reheating.
@@ III Reheating Model with Nonlinear Oscillations
then the leading order form of the potential will have a quartic form when $\phi$ is near its minimum. [^4] Note that one has to translate the field $\phi\to\phi-f$, in order to obtain the first term in equation [8]). Note that the potential (8) represents the simplest model of this type. In particular, the interaction term is linear in $\phi$ and quadratic in $\chi$, which results in a linear equation of motion of the reheating field and requires that the coupling field $\sigma$ is sufficiently small (so that the coupling term does not interfere with the oscillations of the inflaton in its quartic potential). In addition to having a different form at earlier epochs, the inflaton potential could have additional terms (e.g., $V\sim m^{2}\phi^{2}$) during reheating, as well as coupling to additional fields. These complications are considered in subsequent sections below. The interaction between $\phi$ and $\chi$ may take other forms as well, and conformal interaction terms are of particular interest considering the previous discussion. We show in Appendix C that such couplings don’t give favorable results to leading order, so we consider the trilinear interaction throughout this exploration of the quartic inflaton potential’s effects on reheating.
@@ III.1 Solutions for the Inflaton Field
The fluctuations of the inflaton field will experience the equation of motion
$\delta\ddot{\phi}_{k}+\left[k_{\phi}^{2}+m_{\phi}^{2}+3\lambda\phi^{2}+2\sigma\chi\phi\right]\delta\phi_{k}=0\ ,$ (16)
eventually fragmenting the inflaton and breaking the mean-field approximation when $\langle\delta\phi^{2}\rangle\sim\bar{\Phi}^{2}$. The growth of the inflaton fluctuations can be treated similarly to the growth of the fluctuations in the reheat field. However, we do not consider such higher-order effects in this paper and the numeric results show that this growth is subdominant to the growth of the reheat field.
@@ IV Parametric Resonance
Figure 3: Left: Sample trajectories in $(A,q)$ space for $m_{\chi}/\omega_{0}=0$ (top row) and $m_{\chi}/\omega_{0}=2$ (bottom row) for trilinear quartic preheating. The black $A=2q$ line roughly separates the unstable region above from the stable region below. Right: corresponding growth rates $\gamma$ per oscillation time along trajectories as a function of scale factor $a$. We can see that a massless reheat field grows indefinitely without nonlinear effects, while a massive reheat field eventually stops growing on its own when the trajectory reaches a more stable area. Colors correspond to comoving wavenumbers $k/\omega_{0}$ chosen to result in the various $A_{0}$ shown, and the initial inflaton amplitude $\sqrt{\lambda}\Phi_{0}/\omega_{0}$ is chosen to result in $q_{0}\approx 50$.
@@ V Inclusion of a Mass Term in the Inflaton Potential
Figure 5: Growth rates $\gamma$ per oscillation time along trajectories as a function of scale factor $a$ corresponding to the trajectories in Figure 4. The solid lines correspond to the mixed case $m\neq 0,\ \lambda\neq 0$, the dashed lines to purely quartic $m=0,\ \lambda\neq 0$, and the dotted lines to purely massive $m=0,\ \lambda\neq 0$. The three columns correspond from left to right to $\lambda\Phi_{0}^{2}/m^{2}=1,10,1000$ in the mixed potential, and the parameters for the massive and quartic cases are picked to match $(A_{0},q_{0})$ of the mixed cases. The colors correspond to different wavenumbers $k_{0}/\omega_{0}$; note that for the massive case, $\omega_{0}=m$, for the quartic case, $\omega_{0}\approx 0.85\sqrt{\lambda}\Phi_{0}$, and for the mixed case, $\omega$ is the function shown in equation (55). We can see that mass of either the inflaton or reheat field result in an eventual decrease in the growth rate.
@@
VI Numerical Results
Finally, we examine our scenario in fully non-linear lattice simulations. To do this, we employ GABE Child et al. (2013), which evolves the classical equations of motion for both the inflaton
$\ddot{\phi}+3H\dot{\phi}-\frac{\nabla^{2}\phi}{a^{2}}=-\frac{\partial V}{\partial\phi}$ (63)
and the matter field,
$\ddot{\chi}+3H\dot{\chi}-\frac{\nabla^{2}\chi}{a^{2}}=-\frac{\partial V}{\partial\chi}$ (64)
on a uniformly expanding grid. The expansion of the grid is driven by the Friedmann constraint,
$H^{2}=\frac{\langle\rho\rangle}{3m_{p}^{2}}$ (65)
where $\langle\rho\rangle$ is the spatial average of
$\rho=\frac{\dot{\phi}^{2}}{2}+\frac{\nabla^{2}\phi}{2a^{2}}+\frac{\dot{\chi}^{2}}{2}+\frac{\nabla^{2}\chi}{2a^{2}}+V(\phi,\chi).$ (66)
[…]
Figure 6: Variance of the matter field, $\sqrt{\langle\chi^{2}\rangle-\langle\chi\rangle^{2}}$, as a function of time for various choices of $\lambda_{\chi}$. The left panels show model (a), the middle panels show model (b) and the right panel shows model (c). The stars indicate times in each simulation during the system is exhibiting the tachyonic instability. All simulations set $\sigma=2.5\times 10^{-11}\,m_{p}$.
Figure 7: The top panels show the ratio of the power spectrum of $\chi$ during the tachyonic resonance phase, $t_{*}$, to the initial slice, $t_{0}$; these spectra are taken at the points labeled with stars in Figure 6. The bottom panels show the ratio of the power spectrum of $\chi$ at the time of the first peak of the variance of $\chi$, $t_{\rm peak}$, to the initial slice. The left panels show model (a), the middle panels show model (b) and the right panel shows model (c). All simulations set $\sigma=2.5\times 10^{-11}\,m_{p}$
It has long been known that studying tachyonic preheating with three-leg interactions with lattice simulations requires adding a self-coupling, $\lambda_{\chi}\chi^{4}$, to the model so that the classical potential is bounded from below Dufaux et al. (2006a). Figure 6 shows how the variance of $\chi$ evolves over the course of a set of simulations with a fixed $\sigma=2.5\times 10^{-11}\,m_{p}$; as the homogeneous mode of $\phi$ changes sign for the first time, the instability in $\chi$ causes a dramatic rise in this quantity.
Importantly, during this first phase, the $\chi$ self-coupling does not have an effect on the dynamics: this can be seen by the overlapping behavior of the variance in Figure 6 as well as by studying the power spectra of $\chi$. The top row of Figure 7 shows the growth of modes by comparing the power spectrum during the tachyonic phase to the power spectrum at the beginning of the simulation. For a wide range of choices of the self-coupling, these ratios are nearly identical. The bottom row of Figure 7 shows the effect that the self-coupling has by looking at the growth of modes until the first peak of the variance of $\chi$; that is, after the tachyonic instability causes the $\chi$ field to enter the nonlinear regime, the self-coupling of $\chi$ has a dramatic effect on the distribution of power in the $\chi$ field.
Figure 8: Variance of the matter field, $\sqrt{\langle\chi^{2}\rangle-\langle\chi\rangle^{2}}$, as a function of time for two choices of the trilinear coupling: $\sigma=5\times 10^{-13}$ (black, solid) and $\sigma=5\times 10^{-12}$ (red, dashed) . The left panels show model (a), the middle panels show model (b) and the right panel shows model (c). The stars on the plots show times at which we compare the power spectra in Figure 9.
Figure 9: Ratio of the power spectrum of $\chi$ during the first resonance phase, $t_{*}$, to the initial slice, $t_{0}$; these spectra are taken at the points labeled with stars in Figure 6. We directly compare the spectra during this initial instability for all three models at two different values of the trilinear coupling: $\sigma=5\times 10^{-13}$ (solid) and $\sigma=5\times 10^{-12}$ (dashed).
In the cases in Figures 6 and 7, the instability is very strong and we see little difference between the three cases. To see the effect of the inflaton potential on the dynamics, we compare the fiducial $\sigma=5\times 10^{-12}\,m_{p}$ to a lower coupling, $\sigma=5\times 10^{-13}\,m_{p}$. Figure 8 shows that, for this lower choice of the coupling, $\sigma$, reheating does not always complete during the first oscillation of the field. Additionally, we see that the instability is now model-dependent.
We can study the initial instability for each of the three models by looking at the amplification of the power spectrum at the time when the instability occurs (labeled with stars in Figure 8) to the initial power spectra in each case [^7] Note that here we directly plot the ratio of the power spectrum of the field; modes that are not amplified are expected to decrease in amplitude due to the expansion of the universe in each case. Alternatively, we could have chosen to scale the fields to the canonically-normalized fields so that unamplified modes remain constant, as in Adshead et al. (2015). Fig. 9 shows that, for this smaller trilinear coupling, we see a wider and stronger amplification in model (a) as compared to models (b) and (c) which show consistent amplification.
VII Conclusion
@@ A.2 Optimization of the Correction Parameter
Appendix B Comparison of growth rate and expansion rate
Reheating is considered efficient when $\gamma>H$, when the growth rate is greater than the expansion rate of the Universe. Therefore it is worth considering the evolution of the ratio $\gamma/H$; we find that this perspective reinforces the advantage of the quartic potential in increasing the production of $\chi$.
Figure 11: Map of $\gamma/H$ for $V(\phi)=\frac{1}{2}m^{2}\phi^{2}$ (left) and $V(\phi)=\frac{1}{4}\lambda\phi^{4}$ (right). The sample trajectories showcase the general reheating behavior, with trajectories on the left converging on narrow resonance and trajectories on the right tending toward broad resonance. These maps reinforce our conclusion that reheating with a quartic potential is much more efficient than with a massive potential, with some accessory dependence on the relative values of $m$, $\sigma$, and $\lambda$.
Figure 11 shows maps of $\gamma/H$ in the $(A,q)$ plane. In the case where $V(\phi)=\frac{1}{2}m^{2}\phi^{2}$, we see that $\gamma/H$ can actually increase with time, except when crossing narrow stability bands, unlike $\gamma$ alone which decreases with time. Even so, except for a possible burst in the bottom left corner of the map, for most of reheating we have $\gamma/H/(m_{\mathrm{pl}}\sigma/m^{2})<1$. All trajectories converge on $\Phi\ll m^{2}/\sigma$, meaning $\gamma/H\to 0$. In the case where $V(\phi)=\frac{1}{4}\lambda\phi^{4}$, trajectories can easily achieve $\gamma/H/(\lambda m_{\mathrm{pl}}/\sigma)\gg 1$, possibly indefinitely. We know $\gamma\sim\sqrt{q}$ for $q\gg 1$, and it’s clear that $\gamma/H$ increases more quickly with time than $\gamma$ itself. In the case of the quartic potential, we can be more concrete since $\gamma\sim\sqrt{q}$ for large $q$, so that $\gamma/H\sim a^{5/2}$. In all cases, a nonzero reheat field mass $m_{\chi}$ suppresses growth eventually.
Considering $\gamma/H$ further emphasizes the effect of the form of $V(\phi)$ on the efficiency of reheating. The difference can be exaggerated or tempered depending on the value of $\sigma$ relative to the other couplings $m^{2}$ and $\lambda$—notice the units on the colorbars in Figure 11 and their opposite dependence on $\sigma$.
Appendix C Exclusion of quartic terms
@@
C.1 Quartic reheat field potential
A term $V(\chi)\supset\frac{1}{4}\lambda_{\chi}\chi^{4}$ adds a strong restorative force that suppresses the exponential increase in the reheat field’s amplitude, even with modest strength $\lambda_{\chi}<1$. The trilinear equation of motion becomes, with the familiar definitions of $A$ and $q$,
@@ C.1 Quartic reheat field potential
The coefficient increases, $A\to A_{\mathrm{eff}}\equiv A+\lambda_{\chi}\chi_{k}^{2}$, which stabilizes the growth in $\chi_{k}$. The corresponding growth rates are shown in Figure 12. Therefore, we exclude significant $\lambda_{\chi}$ from our analyses to hasten the reheating process. Nevertheless, an appropriate quartic reheat potential guarantees bounding of the potential Dufaux et al. (2006b), a feature we employed to regularize our numerical results in Section VI with $\lambda_{\chi}\ll 1$.
@@
C.2 Forced oscillator solution for first quartic coupling
@@
C.3 Subdominant growth from second quartic coupling